Calculate the theoretical pH of a solution when started with 10 mL of .2 M acetic acid and is diluted to 50 mL with distilled water.
Answer #1

concentration of acetic acid = 10 x 0.2 / 50

                                            = 0.04 M

Ka of acetic acid = 1.8 x 10^-5

[H+] = sqrt (Ka x C)

       = sqrt (1.8 x 10^-5 x 0.04)

       = 8.48 x 10^-4 M

pH = -log [H+] = -log (8.48 x 10^-4)

pH = 3.07